OpenStudy (anonymous):

I need to solve the following problem, 3x over x-2 = 3x+5 over x-1, can anyone help me?

OpenStudy (anonymous):

In this case, you can cross multiply (or multiply both sides by (x-2)(x-1)). When you do that, you get 3x^2 - 3x = 3x^2 -x -10 When you combine like terms, you get 10=2x, so x=5

OpenStudy (anonymous):

Guess you mean: \[\frac{3x}{x-2}=\frac{3x+5}{x-1}\] so you need to say, it's just solvable if x !=2 and x!=1 By cross multiplying u get: \[3x(x-1) = (3x+5)(x-2)\] \[3x^2 -3x = 3x^2 +5x -6x -10 \] \[-3x = -1x -10\] \[-2x = -10\] \[x = 5\]

OpenStudy (anonymous):

Thank you.